ROTATIONAL SPECTRA OF GLYCINE N-METHYL DERIVATIVES: N,N-DIMETHYLGLYCINE AND SARCOSINE

EMILIO J. COCINERO, PABLO VILLANUEVA, M. EUGENIA SANZ, ALBERTO LESARRI, \underline{JUAN C. LÓPEZ}, and JOSÉ L. ALONSO, Grupo de Espectroscopía Molecular (GEM), Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, 47005 Valladolid, Spain.

The conformational behaviour of N,N-dimethylglycine (m.p. 180 ^{∘}C) and sarcosine (m.p. 208 ^{∘}C) has been studied for the first time in gas phase as an extension of our recent work on natural amino acids^{a}. The solid amino acids were vaporized using laser ablation (LA) from a Q-switched Nd:YAG laser, and probed spectroscopically in a supersonic jet using molecular beam Fourier transform microwave spectroscopy (MB-FTMW)^{b}. The rotational spectrum revealed the presence of three conformers of neutral N,N-dimethylglycine in the jet. A bifurcated methyl-to-carbonyl (C–H⋯O=C) weak intramolecular hydrogen bond has been proposed to stabilize the most stable conformer of C_{s} symmetry. The second conformer of C_{1} symmetry presents an intramolecular hydrogen bond N⋯O–H analogue to glycine II. Substitution and effective structures have been derived for this conformer from the rotational data of seven isotopomers. A third conformer presents a cis-carboxylic functional group and C_{1} symmetry. Two conformers were observed for neutral sarcosine. The most stable conformer is stabilized by an intramolecular hydrogen bond N–H⋯O=C, analogue to the interaction observed in the most stable conformer of glycine I. The second stable conformer exhibits an intramolecular hydrogen bond N⋯O–H similar to glycine II. The observed conformers in both molecules are consistent with previous experiments using matrix-isolation FT-IR^{c}.

aA. Lesarri, S. Mata, E. J. Cocinero, S. Blanco, J. C. López, J. L. Alonso, Angew. Chem. Int. Ed., 41 (2002) 4673; A. Lesarri, E. J. Cocinero, J. C. López, J. L. Alonso, Angew. Chem. Int. Ed., 43 (2004) 605; S. Blanco, A. Lesarri, J. C. López, J. L. Alonso, J. Am. Chem. Soc., 126 ( 2004) 11675; A. Lesarri, E. J. Cocinero, J. C. López, J. L. Alonso, J. Am. Chem. Soc., 127 (2005) 2572.

bA. Lesarri, S. Mata, J. C. López and J. L. Alonso, Rev. Sci. Instrum., 74 (2003) 4799.

cA. Gómez-Zavaglia, I. D. Reva, R. Fausto, Phys. Chem. Chem. Phys. 5 (2003) 41; A. Gómez-Zavaglia, R. Fausto, Vib. Spectr., 33 (2003) 105.